OpenStudy (konradzuse):

Use Theorem 4.2.1 to determine which of the following are subspaces of P3

OpenStudy (konradzuse):

(a) All polynomials a0 + a1x + a2x^2 + a3x^3 for which a0 = 0 (b) All polynomials a0 + a1x + a2x^2 + a3x^3 for ewhich a0 + a1 + a2 + a3 = 0 (c) All polynomials a0 + a1x + a2x^2 + a3x^3 of the form in which a0 a1 a2 and a3 are integers (d) All polynomials of the form a0 + a1x , where a0 and a1 are real numbers.

OpenStudy (konradzuse):

I'm confused because it says that W is a subspace of V for both addition and multiplication.

OpenStudy (konradzuse):

It says the answers are A, B, and D, but why not C?

OpenStudy (konradzuse):

@mahmit2012 , @Hero , @jim_thompson5910 any good at Linear?

OpenStudy (anonymous):